OpenStudy (anonymous):

Chad can paint a room in 2 hours. Cassie can paint the room in 3 hours. How long would it take them to paint the room working together?

OpenStudy (anonymous):

6/5 hours

OpenStudy (anonymous):

Work done by Chad and Cassie in 1 hours... \[1/2 + 1/3=5/6\] Thus, total time taken \[6/5\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } + \frac{ 1 }{ 3 }\] \[\frac{ 3+2 }{ 6 }\] Take it's reciprocal : 6/5

OpenStudy (primeralph):

6 years. They'll spend most time arguing lol